OpenStudy (anonymous):

x intercept for the equation 5x-2y=-4?

OpenStudy (anonymous):

-2y=-4-5x y=(-4-5x)/(-2) y=2 + 5/2x so the inter. is 2

OpenStudy (anonymous):

sorry there is a mistake in the last line. Just forget it 0=2+5/2x x=-2*2/5